# Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation

### v1

@@ 4.3. Cross-expert Covariance Loss
Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions. However, they aimed to decorrelate the dimensions within each representation rather than disentangle among representations of all experts. To learn diverse user interests at the expert level, we propose the cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
(6) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p>q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
@@
(7) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov}$
@@ 4.4. Prior Informed Element-wise Gating Mechanism
CovLoss aims to disentangle different representations by minimizing the covariance between dimensions among different experts. However, the existing gating structure hinders the effectiveness of CovLoss. MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020)’s gating mechanism assigns an identical weight to all dimensions within each expert; the dimensions within each expert are mutually constrained. PEPNet (Chang et al., 2023)’s proposed Gate Neaural Unit operates at the dimension level as well. However, it is designed to scale each dimension separately within the representation but not to disentangle these representations. Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig. 3:
@@
Moreover, we are also concerned about the parameter burden of ME structure. In the original ME-MMoE structure, the number of embeddings is tied to the number of domains. This is because the gating is performed using $g^{s}_{k}=Softmax(e_{k}^{s}W_{g}^{s})$, which means the embedding $e_{k}^{s}$ is still related to a $s$-th gating. Thus, when it is deployed in situations with a large amount of domains, the scale of parameters becomes unmanageable. To address this issue, we proposed Prior Informed Gating (PG), which involves an independent set of domain-unrelated prior embeddings $r_{k}$ to control the gating. $r_{k}$ takes the prior features as input, denoted as $x^{prior}$. For a given $k$-th sample and $f$-th field, the looked-up prior embedding is $e_{k,f}^{prior}$. Thus, the input of each gating is:
@@
Furthermore, PG can be understood as an attention (Vaswani et al., 2017)-like mechanism, comprising query, key, and value components corresponding to Prior embedding $r_{k}$, gating weight matrix $W_{g}^{s}$ of $s$-th domain, and expert outputs $\mathbf{O}_{k}$, respectively. Its significance lies in weighting the learnt user interests through the interaction of user, item, and domain information provided by query and key.
4.5. Understanding the CovLoss
CovLoss decorates the representations of different experts in an element-wise fashion. Beyond statistic understanding, CovLoss also directly addresses the core objective of cross-domain modeling: enhance the diversity of representations.
For simplification, we consider the model involving only two experts, outputting $\mathbf{x}$ and $\mathbf{y}$ drawn from two high-dimensional distributions $\mathcal{X}$ and $\mathcal{Y}$, respectively, and the $i$-th element of $\mathbf{x}$ and $\mathbf{y}$ are $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, respectively. Following (Shwartz-Ziv et al., 2024), we assume the randomness comes from the input of the network.
To maximize the diversity of the outputs from both experts, the objective for each pair of samples can be expressed as the expectation of l2-norm difference:
(11) $\mathrm{E}[(||\mathbf{x}||_{2}$ $-||\mathbf{y}||_{2})^{2}]=\mathrm{E}[||\mathbf{x}||_{2}^{2}+||\mathbf{y}||_{2}^{2}-2||\mathbf{x}||_{2}||\mathbf{y}||_{2}]$
$=\sum_{i=1}^{d}[\mu_{xi}^{2}+\sigma_{xi}^{2}]+\sum_{j=1}^{d}[\mu_{yj}^{2}+\sigma_{yj}^{2}]-2\mathrm{E}[\sqrt{\sum_{j\geq i}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{d}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sum_{i=1}^{d}\sum_{j=1}^{d}\sqrt{(x_{i}y_{j})^{2}}]=-2\sum_{j\geq i}\mathrm{E}[x_{i}y_{j}]$
$=-2\sum_{i=1}^{d}\sum_{j=1}^{d}[\mu_{xi}\mu_{y_{j}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}cov(x_{i},y_{j})}],$
where $x_{i},y_{j}\geq 0$ because the outputs from experts have been activated by ReLU before. According to the above Eq., minimizing $cov(x_{i},y_{j})$ can maximize the lower bound of diversity objective [^3] Although $\mu_{xi}\mu_{y_{j}}+cov(x_{i},y_{j})$ simultaneously incorporates both the first and second moments, they are nearly independent of each other in optimization.
@@ 5. Experiment
  • RQ1: How does the Crocodile perform compared with state-of-the-art recommenders in public datasets?
  • RQ2: Can Crocodile capture diverse user interests without introducing dimensional collapse issue?
@@
  • RQ4: How sensitive is Crocodile to critical hyper-parameters and model capacity?
We employed two metrics to measure the model performance: AUC and gAUC. AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
@@ 5.1. Datasets
In this study, to examine the performance of the model in realistic recommendation, we carefully selected the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets for our experiments. For Kuairand1k, we chose five domains with top-5 data volumes. For AliCCP, we selected all three domains. Tab. 1 is the statistics for the two datasets after low-frequency filtering. We removed features with fewer than 10 exposures and replaced them with default values.
@@
#Sample #User #Item #Impr in Domain
Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148 2.4M(S0) / 7.8M(S1) / 0.4M(S2) / 0.9M(S4) / 0.2M(S6)
AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 32M(S1) / 52M(S2) / 0.6M(S3)
@@ 5.2. Baselines
To establish a performance benchmark for comparison, we will select comparative methods from both MTL and MDL for multi-domain recommendation. We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions), and STEM (Su et al., 2023). Besides, we choose representative MDL methods, including STAR (Sheng et al., 2021), PEPNet (Chang et al., 2023), HiNet (Zhou et al., 2023), and AdaSparse (Yang et al., 2022).
5.3. Performance Evaluation (RQ1 and RQ2)
@@
Our proposed Crocodile abrains the best overall AUC and gAUC in two datasets. In the Kuairand1k dataset, Crocodile outperforms the second-best method significantly by 0.09% and 0.19% in terms of overall AUC and gAUC, respectively. In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.04% and gAUC of 0.07% than the second-best method.
Crocodile achieves larger AUC and gAUC almost for almost all domains in two datasets significantly compared to single domain method, with p-value¡0.05. Such positive transfer can be attributed to Crocodile’s better usage of multi-domain information and multi-embedding to enhance performance in each domain, without raising dimensional collapse.
Figure 4. Comparison of the singular value spectrum $log(\sigma)$ of item ID and user ID embeddings in Kuairand1k dataset. We reported $log(\sigma)$ of ME-PLE and STEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
5.3.2. Comparison with MTL for MDR methods.
Multi-task learning (MTL) methods for Multi-Domain Recommendation (MDR) can be primarily categorized into single-embedding (SE) and multi-embedding (ME) paradigms (Su et al., 2023). We observed that some methods allocating domain-specific embedding underperform compared to SE methods, despite the fact they can capture diverse user interests in Fig. 5(a). For instance, in the comparison between PLE and ME-PLE on the Kuairand1k dataset in, Tab. 2, PLE achieved an overall AUC of $0.78565$ and gAUC of $0.66093$. In contrast, ME-PLE’s AUC and gAUC were $0.78499$ and $0.66075$, respectively, both lower than those of PLE.
To investigate the underlying reason, we conducted a domain-by-domain performance comparison and found that ME-PLE exhibited notably poorer results in the smallest domain, S6, with a negative transfer in AUC of -0.4% compared to the single domain. A similar negative transfer was also observed in STEM. We analyzed their embeddings via singular spectrum analysis (Guo et al., 2023b) and show the results in Fig. 4. We discovered that ME-PLE and STEM, in comparison to other methods, exhibited a clear dominance of top factors with much higher singular value over tail factors in the S6 embeddings, indicating the dimensional collapse issue.
In contrast, Crocodile’s embeddings exhibit a more balanced importance of singular values, thereby better supporting the high-dimensional representational space. Consequently, within domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses STEM by 0.37% in AUC and 0.98% in gAUC. Even though STEM learns the most diverse user interest, it still fails to surpass Crocodile due to dimensional collapse.
@@
5.3.3. Comparison with MDL methods.
The Multi-domain Learning (MDL) approaches typically evolve on a trend of progressively allocating domain-personalized components from top to bottom to capture more diverse user interest.
To empirically support this, we analyze the Diversity Index (DI) of both MDL and MTL methods in Fig. 5(a). Firstly, along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and STEM (Su et al., 2023) reveals an increased trend. Secondly, methods with an ME structure typically exhibit higher DI than their SE versions, e.g., ME-PLE has higher DI than PLE.
However, as discussed in Section 5.3.2, the allocation of domain-specific embeddings can lead to the side effect of dimensional collapse, which in turn may result in a negative transfer. Consequently, we are curious as to whether methods that do not employ domain-specific embeddings can achieve favorable outcomes.
For example, HiNet (Zhou et al., 2023) only allocates domain-specific components at the expert level and displayed the best ability to prevent dimensional collapse in Fig. 4. It thus got the best AUC of 0.76451 in domain S6, in Kuairand1k. However, HiNet still falls in suboptimal performance in overall AUC and gAUC in both datasets compared to Crocodile. This is because of its low ability to capture diverse user interests in  5(a).
Conversely, Crocodile achieves the second-best DI, only slightly trailing behind STEM, which employs domain-specific embeddings but faces the issue of dimensional collapse. Consequently, Crocodile attains the best overall AUC and gAUC.
@@
Through comparison with MTL and MDL approaches, we found that Crocodile captures the diverse interests of users (Fig. 5(b)) without introducing the issue of dimensional collapse (Fig. 4), thereby achieving the best overall AUC and gAUC scores in Kuairand1k and AliCCP.
Thus, can other disentangled losses also achieve favorable diversity and corresponding positive outcomes compared to those without them? We compare CovLoss with several other disentangling losses, including Dot Product (Dot) (Wu et al., 2021), Cosine Similarity (Cos) (Liu et al., 2021), Distance Correlation (dCorr) (Liu et al., 2022a) and OLE (Lezama et al., 2018) (F-norm form).
We first analyzed the ability to capture user diverse interests in Fig. 5(b). We observed that all disentangled losses could achieve higher DI than ME-MMoE-PeG, where ME-MMoE-PEG is the basic structure without any disentangled loss compared to others.
Table 4. Comparison of different disentanglement losses with the base model on overall AUC and gAUC in Kuairand1k. * indicates that the performance difference against base is significant at p-value¡0.05.
@@
0.78683 0.07% 0.66373 0.17%
Among them, we found that dCorr (Liu et al., 2022a) loss achieves the smallest improvement of only 0.01% and 0.01% for AUC and gAUC lift, respectively in Tab. 4. This is coherent with its related poor diversity compared to other disentangled losses in Fig. 5(b). In contrast, ultilizing CovLoss achieves the best AUC and gAUC with the lift of 0.07% and 0.17% compared to the Base model. This evidence supports that the ability to learn diverse interests is important to improve performance.
Furthermore, we are interested in exploring whether the transfer of other types of loss functions, e.g., load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), could yield performance enhancements. As shown in Tab. 4, the Importance loss (Shazeer et al., 2017) resulted in a limited improvement, whereas Trans5, one of the effective versions of equivalent transformation losses in (Chen et al., 2023), had a detrimental effect. This can be attributed to that these loss functions do not explicitly promote diversity in learning. Instead, they focus on balancing the contributions of various experts or enhancing the robustness of representations, which does not meet the expected outcomes.

### v2

@@ 3.3. Cross-expert Covariance Loss
In order to make the output of experts uncorrelated so as to capture the user’s diverse or sometimes even conflicting interests on items, inspired by (Bardes et al., 2021; Yu et al., 2020), we propose a novel cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
(7) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p\geq q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
@@
(8) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov},$
@@ 3.4. Prior Informed Element-wise Gating Mechanism
CovLoss aims to disentangle different representations by minimizing the covariance between dimensions among different experts. However, the existing gating structure limits CovLoss’s effectiveness. The gating mechanisms in MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) assign identical weights to all dimensions within each expert, causing mutual constraints among dimensions. PEPNet’s (Chang et al., 2023) Gate Neural Unit operates at the dimension level, but it is designed to scale each dimension separately within the representation rather than to disentangle these representations.
Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig. 3:
@@
However, when the original ME is deployed in situations with a large number of domains, the scale of parameters becomes unmanageable, because the sets of embeddings are still related to the number of domains. To address this issue, we proposed Prior Informed Gating (PG), which involves an independent set of domain-unrelated prior embeddings $r_{k}$ to control the gating. $r_{k}$ takes the prior features as input, denoted as $x^{prior}$. For a given $k$-th sample and $f$-th field, the prior embedding is $e_{k,f}^{prior}$ and the input of each gating is:
@@
3.5. Discussion
The Covariance Loss
Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions. Furthermore, some (Hua et al., 2021) claimed that feature de-correlation by standardizing the covariance matrix can help representations overcome dimensional collapse. These efforts have provided evidence to demonstrate the interpretability and efficacy of covariance. However, they aimed to de-correlate the dimensions within each representation rather than disentangle among representations of all experts. Besides, this loss shares the same spirit with the Maximal Coding Rate Reduction Principle (Chan et al., 2022; Yu et al., 2020) to de-correlate representations from different clusters (experts in our model).
Figure 4. Architecture comparison of different MDL models with multiple embeddings.
Connection to Existing Mulit-Embedding MTL or MDL Methods
The proposed Crocodile differs from those existing multi-embedding-based methods in the following aspects: 1. Different from ME-PLE (Tang et al., 2020; Su et al., 2023) and SDEM (Su et al., 2023), which have domain-specific embeddings, all the embeddings and experts in Crocodile and ME-MMoE are shared across domains. Such design alleviates the insufficient learning of those embeddings corresponding to the small domains as illustrated in Sec. 2.2. 2. The embeddings and experts in Crocodile are disentangled with the Covariance Loss, making it able to preserve the distinctiveness of each domain, as illustrated in Sec. 2.1. Fig. 4 shows an architecture comparison of these multi-embedding MDL methods.
@@ 4. Experiment
  • RQ1: Can Crocodile outperform state-of-the-art methods on public datasets? Can this be attributed to maintaining domain distinctiveness while enabling sufficient representation learning?
  • RQ2: Can CovLoss effectively disentangle the expert outputs?
@@
  • RQ4: How sensitive is Crocodile to hyper-parameters and model capacity?
  • RQ5: How does Crocodile perform on a large-scale online recommendation system?
The two public datasets correspond to video and product recommendations, respectively. The online revenue of these scenarios is often related to the ranking performance, ensuring that higher-scoring items better match user interests. Hence, we employed two metrics to measure the model performance: AUC and gAUC. AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
@@ 4.1. Datasets
In this study, to examine the performance of the model in realistic recommendation, we carefully selected the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets for our experiments. For Kuairand1k, we chose five domains with top-5 data volumes. Kuairand1k has 30 user-side and 62 item-side features. For AliCCP, we selected all three domains. AliCCP has 12 user-side, 4 item-side, and 4 user-item interaction features. Tab. 1 is the statistics for the two datasets after low-frequency filtering. We removed features with fewer than 10 exposures and replaced them with default values.
Both of these datasets exhibit issues of extreme data imbalance. For instance, in the Kuairand1k dataset, there is a 12-fold difference between the data in S0 and S6. Similarly, in the AliCCP dataset, S2 has 86 times larger data than S3.
@@
#Sample #User #Item Domain (#Impr)
Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148
S0(2.4M)/S1(7.8M)/S2(0.4M)
S4(0.9M)/ S6(0.2M)
AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 S1(32M) / S2(52M) / S3(0.6M)
@@ 4.2. Baselines
To establish a performance benchmark for comparison, we will select comparative methods from both MTL and MDL for multi-domain recommendation. We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions). Inspired by STEM (Su et al., 2023), we also proposed the SDEM method with Shared and Domain-specific EMbeddings. Besides, we choose representative MDL methods, including STAR (Sheng et al., 2021), PEPNet (Chang et al., 2023), HiNet (Zhou et al., 2023), and AdaSparse (Yang et al., 2022).
4.3. Performance Evaluation (RQ1)
@@
Our proposed Crocodile achieves the best overall AUC and gAUC in two datasets. In the Kuairand1k dataset, Crocodile outperforms the second-best method significantly by 0.09% and 0.19% in terms of overall AUC and gAUC, respectively. In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.138% and gAUC of 0.12% than the second-best method. Usually, a 0.1% AUC (gAUC) lift is regarded as a huge improvement in recommendation systems (Zhu et al., 2022).
Crocodile achieves larger AUC and gAUC almost for almost all domains in two datasets significantly compared to single domain method, with p-value¡0.05. Such positive transfer can be attributed to Crocodile’s preserving domain distinctive and sufficiently representation learning.
Figure 5. Comparison of the singular value spectrum $log(\sigma)$ of item ID and user ID embeddings in Kuairand1k dataset. We reported $log(\sigma)$ of ME-PLE and SDEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
4.3.2. MDL methods with domain-specific embeddings.
The Mulit-Embedding PLE (ME-PLE) and SDEM are two new variants with domain-specific embeddings, inspired by task-specific embeddings in multi-task learning. However, ME-PLE’s overall AUC and gAUC are lower by 0.07% and 0.02% compared to PLE, respectively. Analyzing each domain, we found the smallest domain experienced the most significant drops in AUC and gAUC, both exceeding 0.3%. A similar negative transfer was also observed in SDEM.
To further investigate the mechanism, we examined the quality of representation learning by observing the imbalance in the distribution of singular values (Guo et al., 2023b; Chi et al., 2022). As shown in Fig. 5, We discovered that ME-PLE and SDEM exhibited a clear dominance of top factors with much higher singular values over tail factors in the S6 embeddings. This means that their embeddings of small domains suffer from insufficient learning. In contrast, Crocodile’s embeddings exhibit a more balanced importance of singular values, thereby better supporting the high-dimensional representational space. Consequently, within the smallest domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses SDEM by 0.37% in AUC and 0.98% in gAUC. Even though SDEM learns the most diverse user interests, it still fails to surpass Crocodile due to insufficient representation learning.
@@
((a)) CovLoss between Experts
((b)) Covariance across all dimensions of all Experts
Figure 7. Heatmaps of CovLoss between Experts (left) and Covariance across all dimensions of all Experts (right) on Kuairand1k. In the comparison methods with the same structure, ”without CovLoss” corresponds to the loss function being only BCE Loss, while ”with CovLoss” corresponds to the loss function as the combination of BCE Loss and CovLoss.
4.3.3. Existing MDL methods.
Multi-domain learning (MDL) methods also exhibit a top-down structural evolution, explicitly capturing domain-specific interests by allocating domain-specific components layer by layer. We expect that the more thorough the design, the more domain-distinctive the learnable parameters, resulting in better performance. We use the IA metric introduced in Sec. 2 to quantify it, as shown in Fig. 6(a). Along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), PEPNet (Chang et al., 2023), and SDEM (Su et al., 2023) reveals an increasing trend, which is consistent with expectations. However, the performance does not follow this order. For example, SDEM has a better DI than HiNet but exhibits lower AUC and gAUC, because performance is also influenced by the sufficiency of representation learning, as discussed in Sec. 2.2.
In contrast, Crocodile maintains domain distinctiveness more effectively while ensuring adequate representation learning. Although HiNet exhibits more balanced singular values compared to Crocodile, it performs poorly in maintaining domain distinction. Consequently, Crocodile achieves higher AUC and gAUC than HiNet by 0.12% and 0.24%, respectively.
@@
In Sec. 4.3.2 and  4.3.3, we found that Crocodile can achieve sufficient representation learning while maintaining domain distinctiveness, resulting in the best performance on two public datasets. This raises the question: can other loss functions achieve similar benefits? We conducted experiments by replacing the loss function in Crocodile, naming the method that retains the structure but uses only BCE Loss as Base.
Table 4. Comparison of different losses with the base model on overall AUC and gAUC in Kuairand1k. * indicates that the performance difference against base is significant at p-value¡0.05.
@@
0.78683* 0.07% 0.66373* 0.17%
Among the loss functions in Tab. 4, dCorr (Liu et al., 2022a) shows the smallest improvement, with only a 0.01% increase in AUC and gAUC, consistent with its poor domain distinctiveness in Fig. 6(b). In contrast, CovLoss achieves the highest AUC and gAUC improvements of 0.07% and 0.17% over the Base model, highlighting the importance of domain distinctiveness for performance enhancement.
Additionally, we explored other loss functions, such as load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), for potential performance gains. As shown in Tab. 4, Importance loss (Shazeer et al., 2017) offered limited improvement, while Trans5, an effective transformation loss from (Chen et al., 2023), had a negative impact. They all fail to surpass CovLoss.
4.4. CovLoss Analysis (RQ2)
We are curious about if the CovLoss truly disentangled the experts as expected. We conducted experiments using the same backbone on Kuairand1k, with and without CovLoss in Fig. 7. We progressively reveal the effectiveness of CovLoss from the course, i.e., expert level, to fine, i.e., dimension level.
Firstly, at the expert level in Fig. 7(a), the method with CovLoss shows significantly lower CovLoss values across each pair of experts compared to the baseline. For example, the CovLoss between the 3-th and 5-th Experts in the method with CovLoss is $3\times 10^{3}$ lower than that without CovLoss.
Secondly, at the dimension level in Fig. 7(b), the dimensional-wise heatmap of the method with CovLoss has much smaller number of high-value points than without CovLoss, shows it also has overall significantly smaller covariance compared to the baseline.
Please note that CovLoss acts as an auxiliary loss, serving as a soft constraint and certain learnable parameters must retain intensity information. Thus, few high-value points are still observed in the method with CovLoss.
Finally, in terms of experimental performance, the method with CovLoss achieved AUC and gAUC significant improvements of 0.07% and 0.17%, respectively, on the Kuairand1k dataset compared to the method without CovLoss, as shown in Tab. 5. These findings demonstrate that CovLoss effectively disentangles the experts and enhances model performance.